# Continuous-time reinforcement learning for optimal switching over multiple regimes

### v1

@@ 6.1 Bounded Regulator Problem
We use a discrete version of (6.1) for $t=0,\Delta t,...,K\Delta t$ with $K=100$ and $\Delta t=T/K$. The value function and policy are approximated by a neural network in the PyTorch framework with the architecture and parameters summarized in Table 1.
@@
Figure 1: (a): Convergence of the training loss for the bounded regulator problem with $\lambda=0.2$. (b): Learned value functions and switching probabilities at $t=0.5$ for $\lambda=0.2$.
@@ 6.2 Put Option Selection Problem
with $S^{A}_{0}=s^{A}\in[0,\infty),S^{B}_{0}=s^{B}\in[0,\infty)$. Here the parameters are set by $(\mu^{A},\sigma^{A})=(0.1,0.2)$ and $(\mu^{B},\sigma^{B})=(0.05,0.1)$, and $W=(W_{t})_{t\in[0,T]}$ is a standard Brownian motion. The risk free rate is $r=0.05$. For any time $t\in[0,T]$, the investor decides a action $I_{t}\in\{0,1,2\}$, which determines the regime in which the investor’s wealth is allocated. Switching between regimes incurs transaction costs given by the matrix:
@@
According to Figure 4, at the beginning of training, the loss exhibits a oscillation, and the convergence is very pronounced. It becomes stable when the number of episodes exceeds 800. Figure 5 shows the allocation of the asset at time $t=0.5$. We can find that, when stock price of $A$ and $B$ large enough, the investor will put all in bank. When stock B has lower price, she tends to hold put $A$; When stock $A$ has lower price, she tends to hold put $B$.
@@
Figure 5: The optimal asset allocation policy at $t=0.5$ as a function of stock prices $S^{A}$ and $S^{B}$.

### v2

@@ 6.1 Bounded Regulator Problem
We use a discrete version of (6.1) for $t=0,\Delta t,...,K\Delta t$ with $T=1$, $K=100$ and $\Delta t=T/K$. The value function and policy are approximated by a neural network in the PyTorch framework with the architecture and parameters summarized in Table 1.
In order to approximate the expectation operator in the theoretical formulation via Monte Carlo methods—that is, to estimate the gradient direction using the average of finite samples so that the update of parameters $\xi$ is not based on a single specific scenario—we employ batch learning to train the neural network parameters $\xi$. This approach not only enhances training stability but also ensures that the learned policy is optimal in an expected sense.
@@
Figure 1: (a): Convergence of the training loss for the bounded regulator problem with $\lambda=0.2$. (b): Learned value functions and switching probabilities at $t=0.5$ and $I_{0.5}=0$ for $\lambda=0.2$, i.e, $P(I_{t+\Delta t}=1|I_{t}=0,X_{t}=x,t=0.5)$.
@@ 6.2 Put Option Selection Problem
with $S^{A}_{0}=s^{A}\in[0,\infty),S^{B}_{0}=s^{B}\in[0,\infty)$. Here the parameters are set by $(\mu^{A},\sigma^{A})=(0.1,0.2)$ and $(\mu^{B},\sigma^{B})=(0.05,0.1)$, and $W=(W_{t})_{t\in[0,T]}$ is a standard Brownian motion. The risk free rate is $r=0.05$. For any time $t\in[0,T]$, the investor decides an action $I_{t}\in\{0,1,2\}$, which determines the regime in which the investor’s wealth is allocated. Switching between regimes incurs transaction costs given by the matrix:
@@
According to Figure 4, at the beginning of training, the loss exhibits a oscillation, and the convergence is very pronounced. It becomes stable when the number of episodes exceeds 800. We then plot in Figure 5 the learnt value functions $(V^{1},V^{2},V^{3})$. Conforming to financial intuition, each value function $V^{i}$ (for $i=1,2,3$) is a decreasing function of both underlying stock prices $s^{A}$ and $s^{B}$.
@@
Figure 5: Left panel: The learnt value functions $s^{A}\to(V^{1},V^{2},V^{3})$ with $(s^{B},t)=(1.0,0.5)$. Right panel: The learnt value functions $s^{B}\to(V^{1},V^{2},V^{3})$ with $(s^{A},t)=(1.0,0.5)$.